OpenStudy (anonymous):

MEDAL Given the following functions f(x) and g(x), solve f over g(−4) and select the correct answer below: f(x) = 4x − 4 g(x) = x − 1 −20 −4 4 one fourth

OpenStudy (anonymous):

f/g (-4) is the function

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

help @Haseeb96 :D?

OpenStudy (haseeb96):

look we will solve this question in this way g(x)= x-1 put x=-4 in the above equation f(x)/g(x)= 4x-4 / x-1 =4(x-1)/x-1 = 4 f(x)/g(-4)= 4 is your correct answer

OpenStudy (anonymous):

4 is my answer?

OpenStudy (anonymous):

4 was right, thank you :)

OpenStudy (haseeb96):

yes